OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 3x – 4y = 8?

OpenStudy (anonymous):

if they're parallel they have the same slope/gradient 3x-4y=8 -4y=8-3x 4y=3x-8 y=3/4x-2 hence slope/gradient is 3/4

OpenStudy (demandandsupply):

As i said, Auxuris is correct...